OpenStudy (anonymous):

if ab=6,c-b=5, 2c=16, what is the value of a+b+c? Please help me :/

OpenStudy (pinksapphire):

a=2 b=3 c=8

OpenStudy (pinksapphire):

2+3+8=13

OpenStudy (lopus):

c=16/2 c-b=5 -b=5-c ab=6

OpenStudy (anonymous):

Omg thank you so much! Which one is the real way to get the solution though?

OpenStudy (lopus):

multiplication A*B=C then A=C/B or B=A/C add A+B=C the A=C-B or B=C-A

OpenStudy (lopus):

sorry you need de answer or know how

OpenStudy (pinksapphire):

I have the answers.

OpenStudy (anonymous):

Thank you I got yours @PinkSapphire

OpenStudy (lopus):

have a mistake multiplication A*B=C then A=C/B or B=C/A add A+B=C the A=C-B or B=C-A